The four particles shown in Fig. connected by rigid rods. The or igin is at the center of the rectangle. If the system rotate in the xy plane about the z axis with an angular speed of 8.00 rad/s, calculate yim) 9.00 kg 2.00kg 04: The moment of inertia of the system about the z axis (A) 143Kg. m2 (B) 65Kg.m2 (C) 78Kg. m2 (D) 52Kg m2 (E) 91Kg.m2 6.00 m x(m) 0 CS: The rotational kinetic energy of the system. ( A) 3.50 × 1oY (B) 4.58 × 103, (C) 1.14 × 10ッ (D) 1.78 x 10J (E)2.57 x 10 200 kg 4.00 kg Q6 A 0.500-kg ball is thrown straight up into the air with an initiai spe ed of 15.0 m/s. Find the momen tuin of The ball at halfway up to its maximum height (A) 5.:30Kg.ms (B)1.06Kg.ms 212Kg.s (D)3.18g.ms-1 (E) 4.24Kg.ms-
Answer #1

Q4)

Iz = Ix + Iy = My^2 + Mx^2

= 2(3^2+2^2) + 3(3^2+2^2) + 2(3^2+2^2) + 4(3^2+2^2)

= 11 x 13 = 143 kgm^2

Q5)

K = 1/2 I w^2

- 143/2 x 8^2 = 4576 J = 4.58 kJ

Q6

h = v^2/2g = 15^2 / 20 = 11.25 m

v = (2gh)^0.5 = (2x10x5.625)^0.5

= 10.61 m/s halfway

p = mv = 0.5x10.61 = 5.3 kgm/s